The length of a rectangle is equal to 3 times it's width. If the perimeter is equal to 240 ft. What is the lenght of the rectangle?
Answer by mangopeeler07(462)

length=3w
width=w

2(3w)+2(w)=240

6w+2w=240
8w=240
w=30
length=3(30)

length=90